Grade 12th passGeneral Physics

A ray of light incident normally on the face of a prism of refractive index= 1.5 with equilateral principal section. The angle between the incident and emergent rays would be:
(1) 90° (2) 180° (3) 120° (4) 60°

Indrani Majumdar
10 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago
angle of deviation in prism is:
S = i + e – A
light incident normally on the face .i.e.
i = 0 and  r1 = 0 r2 = 60 degree.
r2 > critical angle .i.e = arcsin(1/1.5) = 41.81 degree. 
S = 0 – 1 + 60 = degree.